The following information is available from the accounting records of Eva Corporation. Fixed costs per period are​ $4800. Sales volume for the last period was ​$19,360​, and variable costs were ​$13,552. Capacity per period is a sales volume of ​$32,000.
(i)what is the total contribution margin

Answer by Theo(13342)

the total contribution margin is the the total sales price minus the total variable cost.

that would make it 19360 - 13552 = 5808.

the total contribution margin rate is (the total sales price minus the total variable cost) divided by the total sales price.

that would make it 5808 / 19360 = .3 = 30%.

at capacity, the total sale price would be 32000.

at 30% total contribution margin rate, the total contribution margin would be .3 * 32000 = 9600.

here's some references on total contribution margin.

https://www.accountingtools.com/articles/2017/5/16/contribution-margin

https://xplaind.com/325438/contribution-margin

https://www.thebalancesmb.com/what-is-the-contribution-margin-393476

there are plenty more references.

just do a search for total contribution margin and they'll pop up.

here's an edample of the search that i did using google.

https://www.google.com/search?q=total+contribution+margin&rlz=1C1CHBF_enUS839US839&oq=total+contribution+margin&aqs=chrome.0.69i59j0l5.10335j1j4&sourceid=chrome&ie=UTF-8